OpenStudy (anonymous):

Solve: 2(x + 1) + 4 = 8

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

2(x+1)+4=8 2x+2+4=8 Distribute 2x+6=8 Add liked terms 2x=2 Divide both sides by 2 x=1

OpenStudy (anonymous):

yes exactly!